OpenStudy (anonymous):

How much energy (in kilojoules) is needed to heat 3.10g of ice from -10.5∘C to 22.5∘C? The heat of fusion of water is 6.01kJ/mol, and the molar heat capacity is 36.6 J/(K⋅mol) for ice and 75.3 J/(K⋅mol) for liquid water.

OpenStudy (aaronq):

you have to find 3 things, then add them all up. 1) heat required to raise the T of ice from -10.5 to 0 2) heat required to melt the ice. use: q=n*H(fusion) ; H(fusion)=6.01kJ/mol 3) heat required to raise the T of the water from 0 to 22.5 for 1) & 3) use: q=m*C*dT where, dT=T(final)-T(initial) q = heat m= mass C=specific heat capacity